OpenStudy (anonymous):

(3x+5)+(2x+1)=5(x-1)+9

OpenStudy (anonymous):

I'm not going to do all your homework for you :P Solution is the same here as before, combine x terms on one side, factor out and divide off the rest.

hero (hero):

5x + 6 = 5x + 4 No Solution

OpenStudy (anonymous):

3x + 5 + 2x + 1 = 5x -5 + 9 3x + 2x + 5 + 1 = 5x - 5 + 9 5x + 5 + 1 = 5x - 5 + 9 5x + 6 = 5x - 5 + 9 6 = 5 + 9 6 = 14 Therefore this equation makes no sense, and has no solution.

OpenStudy (anonymous):

I had a type, 6 = -5 + 9 in the second to last line and 6 = 4 in the last line